# Laura Barros        Aiko Schmeißer

### v1

@@ 7.2 Local-level evidence
(1) (2) (3) (4) Outcome: $\Delta$ Average years of education of … Councilor candidates Elected councilors [A] OLS $\Delta$ Employment rate (ppt) -0.047∗∗∗ -0.043∗∗∗ -0.038∗∗∗ -0.035∗∗∗ (0.005) (0.004) (0.008) (0.007) Observations 16,651 16,651 16,651 16,651 [B] 2SLS (Bartik IV) $\Delta$ Employment rate (ppt) -0.165∗∗∗ -0.157∗∗∗ -0.144∗∗∗ -0.135∗∗∗ (0.018) (0.016) (0.022) (0.022) KP F-stat 124.4 194.4 124.4 194.4 Observations 16,651 16,651 16,651 16,651 Fixed effects Year State $\times$ year Year State $\times$ year Note: The table presents OLS and 2SLS results from model (5) for the effect of formal employment rates on the average education of local councilors. The sample contains all municipal elections in 2004, 2008, 2012, and 2016. The dependent variable measures the change across subsequent elections in the average education of all individuals running (columns (1) and (2)) and being elected (columns (3) and (4)) for local council. The explanatory variable is the percentage point change in the municipal formal employment-to-population rate. For the 2SLS results, changes in employment rates are instrumented with Bartik instruments that are constructed by interacting baseline differences in municipalities’ industry composition with national industry employment growth rates (see equation (6)). All regressions are weighted by municipal population. Standard errors clustered at the municipality level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
Table 5 presents the results of OLS and 2SLS regressions for model (5). In all specifications, we see a significantly negative relation between local employment rates and the average education of those who run for local council. The Bartik IV regression implies that a 1 percentage point lower employment rate increases candidates’ education by 0.17 years. This result is consistent with our findings at the individual level, which indicate that job loss has larger positive effects on candidacies among better-educated individuals. Importantly, Table 5 also shows that local labor demand shocks affect not only the selection of those who run for office but also those who are ultimately elected to office. A 1 percentage point lower employment rate is estimated to increase the education of elected councilors by 0.14 years. Together, these results highlight that negative economic shocks can have positive effects on the selection of local politicians.
@@ 8 Conclusion
The quality of democratic institutions largely depends on individuals’ willingness to engage in politics. While political careers likely compete with careers in the private sector, we lack causal evidence on how changes in labor market opportunities affect the trade-offs faced by individuals when deciding to enter politics. Combining rich information on the universe of party members and local candidates with matched employer-employee data covering all formal workers in Brazil, we study the effects of mass layoffs on individuals’ decisions to start a political career and the consequences of economic shocks for the selection of politicians.
Our results show that layoffs significantly increase individuals’ likelihood of joining a political party or running for office in local elections. These positive effects persist over time, remaining statistically significant up to 12 years after the layoff. Furthermore, they are prevalent across a wide range of individual, regional, and party characteristics and extend to alternative forms of political participation, including running for mayor, donating to political campaigns, and participating in political campaign work.
@@
Several of our results suggest that the increase in political entry can be rationalized by an opportunity-cost mechanism, where job loss reduces individuals’ alternative options in the private sector and increases their availability of time resources. In particular, we observe larger increases in candidacies among individuals with higher financial incentives – measured either as wages in the public office or the predicted income loss due to the layoff. Additionally, we find that eligibility for unemployment insurance benefits, by increasing the time outside of employment, positively affects political entry.
@@
In terms of the implications of layoffs for the selection of political entrants, we document that treated individuals who become party members or candidates after their layoff had, on average, higher wages prior to layoff, higher education levels, and more political experience, as compared to control candidates. This positive selection is also evident at the aggregate level, where we find a positive effect of municipal labor demand shocks on the education of individuals running for and being elected to local councils. Taken together, our paper provides novel evidence that layoffs increase the likelihood that highly capable individuals enter politics. This suggests that political selection can be counter-cyclical: negative economic shocks lead to an increase in politician quality.
@@
While our results emphasize how economic trade-offs can influence who joins politics, they also raise important questions about how private-sector opportunities shape the behavior of politicians once they attain power. In particular, do politicians’ economic circumstances influence their performance in office, including their legislative productivity, their propensity for personal rent extraction, and their ability to effectively provide local public goods? Furthermore, how does experiencing a job loss impact the policy priorities set by politicians and the representation of voters’ preferences in the policy-making process? These remain unresolved questions that are crucial for understanding the broader consequences of economic shocks for the supply side of the political system.
Appendix
Appendix A Appendix to Sections 2 to 4
@@
2004 2008 2012 2016 2020 All Total donations Mean 3,715.55 4,861.25 5,649.81 3,427.54 3,467.26 4,200.86 P10 202.69 211.58 212.38 230.70 214.83 213.61 P25 520.92 617.11 613.78 534.02 519.33 552.17 P50 1,403.30 1,833.70 2,123.80 1,468.55 1,327.48 1,586.85 P75 3,506.12 4,752.17 5,932.49 3,741.29 3,276.06 4,169.63 P90 7,994.55 10,655.85 13,199.48 7,435.77 7,450.96 9,300.73 Donations from own resources Share$>$0 0.61 0.68 0.55 0.56 0.46 0.56 Mean 1,632.70 1,718.96 1,704.98 1,225.80 624.07 1,313.14 P10 0.00 0.00 0.00 0.00 0.00 0.00 P25 0.00 0.00 0.00 0.00 0.00 0.00 P50 318.93 439.03 196.81 213.61 0.00 168.13 P75 1,811.53 1,819.59 1,792.28 1,281.64 681.86 1,275.73 P90 4,297.07 4,557.80 4,955.54 3,226.50 1,494.49 3,539.67 N 202,534 284,031 359,753 395,179 415,691 1,657,188 Note: The table shows summary statistics for campaign donations received by candidates for local council in the elections 2000 to 2020. We report total donation values and the value of donations that come from candidates’ own resources. All values are in BRL (CPI 2018) and are winsorized at the 1% level. The data is obtained from the Tribunal Superior Eleitoral.
@@
Table A.3: Summary statistics: party characteristics
@@
Appendix B Appendix to Section 5
Figure B.1: Labor market outcomes
A. Mass layoff B. Job separation Refer to caption Refer to caption C. Employment D. Labor earnings Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on the likelihood of being displaced in a mass layoff (Panel A), the likelihood of having any job separation (Panel B), the likelihood of being full-time employed at the end of the year (Panel C), and yearly labor earnings (Panel D). $N=24,549,564$ individual-year observations. The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group in the post-treatment period are reported.
@@ Appendix B Appendix to Section 5
Figure B.8: Heterogeneity by municipality characteristics
Table B.4: Councilor wage caps

### v3

@@ 6.2 Local-level evidence
(1) (2) (3) (4) Outcome: $\Delta$ Average years of education of … Councilor candidates Elected councilors [A] OLS $\Delta$ Employment rate (ppt) -0.035∗∗∗ -0.029∗∗∗ -$0∗∗$027∗∗∗ (0.003) (0.004) (0.006) (0.007) Observations 16,650 16,650 16,650 16,650 [B] 2SLS (Bartik IV) $\Delta$ Employment rate (ppt) -0.160∗∗∗ -0.150∗∗∗ -0.140∗∗∗ -0.122∗∗∗ (0.014) (0.016) (0.022) (0.025) KP F-stat 301.7 250.3 301.7 250.3 Observations 16,650 16,650 16,650 16,650 Fixed effects Year State $\times$ year Year State $\times$ year Note: The table presents OLS and 2SLS results from model (5) for the effect of formal employment rates on the average education of local councilors. The sample contains all municipal elections in 2004, 2008, 2012, and 2016. The dependent variable measures the change across subsequent elections in the average years of education of all individuals running (columns (1) and (2)) and being elected (columns (3) and (4)) for local council. The explanatory variable is the percentage point change in the municipal formal employment-to-population rate. For the 2SLS results, changes in employment rates are instrumented with Bartik instruments that are constructed by interacting baseline differences in municipalities’ industry composition with national industry employment growth rates (see equation (6)). All regressions are weighted by municipal population. Standard errors clustered at the municipality level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
Table 4 presents the results of OLS and 2SLS regressions for model (5). In all specifications, we see a significantly negative relation between local employment rates and the average education of those who run for local council. The Bartik IV regression implies that a 1 percentage point lower employment rate increases candidates’ education by 0.16 years. This result is consistent with our findings at the individual level, which indicate that job loss has larger positive effects on candidacies among better-educated individuals. Importantly, Table 4 also shows that local labor demand shocks affect not only the selection of those who run for office but also those who are ultimately elected to office. A 1 percentage point lower employment rate is estimated to increase the education of elected councilors by 0.14 years.
Appendix Tables D.3 and D.4 further show that negative local labor demands shocks can also improve political representation. Consistent with our individual-level analysis of compliers, we find that lower employment rates lead to an increase in the share of female and younger individuals who run for and are elected to councilor office. Together, these results highlight that negative economic shocks can enhance the competence of local politicians without compromising their socio-demographic representation.
@@ 7 Conclusion
The quality of democratic institutions largely depends on individuals’ willingness to engage in politics. While political careers plausibly compete with careers in the private sector, we lack causal evidence on how changes in labor market opportunities affect the trade-offs faced by individuals when deciding to enter politics. Combining rich information on the universe of party members and local candidates with matched employer-employee data covering all formal workers in Brazil, we study the effects of mass layoffs on individuals’ decisions to start a political career and the consequences of economic shocks for the selection of politicians.
@@
Our results show that layoffs significantly increase individuals’ likelihood of joining a political party or running for office in local elections. These positive effects persist over time, are prevalent across a wide range of individual, regional, and party characteristics, and extend to alternative forms of political participation. Several of our results suggest that the increase in political entry can be rationalized by an opportunity-cost mechanism, whereby job loss reduces individuals’ outside options in the private sector and increases their available time resources. In particular, we observe larger increases in candidacies among individuals with higher financial incentives – measured either by the predicted income loss from the layoff or the wage received in public office – and among those with more time resources which result from UI benefit incentives.
@@
In terms of the implications of layoffs for the selection of political entrants, we document that treated individuals who become party members or candidates because of the layoff are positively selected across a number of competence measures. This positive selection is also evident at the aggregate level, where we find a positive effect of municipal labor demand shocks on the education of individuals running for and being elected to local councils. Taken together, our paper provides novel evidence that layoffs increase the likelihood that highly capable individuals enter politics. This suggests that political selection can be counter-cyclical: negative economic shocks may lead to an increase in politician quality.
@@
While our results emphasize how economic trade-offs can influence who joins politics, they also raise important questions about how private-sector opportunities shape the behavior of politicians once they attain power. In particular, do politicians’ economic circumstances influence their performance in office, including legislative productivity, propensity for personal rent extraction, and ability to effectively provide public goods? Furthermore, how does experiencing a job loss impact the policy priorities set by politicians and the representation of voters’ preferences in policy-making? These remain unresolved questions that are crucial for understanding the broader consequences of economic shocks for the supply side of the political system.
Supplemental Appendix
Appendix A Appendix to Sections 2 to 3
Figure A.1: Employment effects of job loss by party membership before the layoff
A. Public-sector employment B. Private-sector employment Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on the likelihood to be formally employed at the end of the year in the public sector (Panel A) and in the private sector (Panel B), separately for individuals who were or were not member of political party before the layoff. The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group across all post-treatment periods are reported.
@@ Appendix A Appendix to Sections 2 to 3
2004 2008 2012 2016 2020 All Total donations Mean 3,715.55 4,861.25 5,649.81 3,427.54 3,467.26 4,200.86 P10 202.69 211.58 212.38 230.70 214.83 213.61 P25 520.92 617.11 613.78 534.02 519.33 552.17 P50 1,403.30 1,833.70 2,123.80 1,468.55 1,327.48 1,586.85 P75 3,506.12 4,752.17 5,932.49 3,741.29 3,276.06 4,169.63 P90 7,994.55 10,655.85 13,199.48 7,435.77 7,450.96 9,300.73 Donations from own resources Share$>$0 0.61 0.68 0.55 0.56 0.46 0.56 Mean 1,632.70 1,718.96 1,704.98 1,225.80 624.07 1,313.14 P10 0.00 0.00 0.00 0.00 0.00 0.00 P25 0.00 0.00 0.00 0.00 0.00 0.00 P50 318.93 439.03 196.81 213.61 0.00 168.13 P75 1,811.53 1,819.59 1,792.28 1,281.64 681.86 1,275.73 P90 4,297.07 4,557.80 4,955.54 3,226.50 1,494.49 3,539.67 Observations 202,534 284,031 359,753 395,179 415,691 1,657,188 Note: The table shows summary statistics for campaign donations received by candidates for local council in the elections 2000 to 2020. We report total donation values and the value of donations that come from candidates’ own resources. All values are in BRL (CPI 2018) and are winsorized at the 1% level. The data is obtained from the Tribunal Superior Eleitoral.
@@
Appendix B Appendix to Section 4
Table B.1: Summary statistics: party characteristics
@@ Appendix B Appendix to Section 4
Figure B.1: Effects of job loss on future separations
A. Mass layoff B. Any job separation Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on the likelihood of being displaced in a mass layoff (Panel A) and of having any job separation (Panel B). $N=24,549,564$ individual-year observations. The graphs include vertical bars that depict 95% confidence intervals based on standard errors clustered at the individual level (too small to be visible). Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group in the post-treatment period are reported.
@@
Figure B.4: Heterogeneity by municipality characteristics
Figure B.5: Labor market effects by education level
A. Labor earnings B. Any employment Refer to caption Refer to caption C. Private-sector employment D. Public-sector employment Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on yearly formal labor earnings (Panel A), the likelihood of being employed in any formal job at the end of the year (Panel B), and the likelihoods of being employed in the private sector (Panel C) and public sector (Panel D), separately for individuals with no degree, elementary school, or middle school ($<$ 12 years), high school (12 years), and university education ($>$ 12 years). The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group across all post-treatment periods are reported.
Table B.4: Councilor wage caps
@@ Appendix C Appendix to Section 5
Appendix D Appendix to Section 6
Table D.1: Summary of AKM estimation
Largest connected set
Share of sample 0.983
Number of worker-year observations 189,819,461
Number of establishments 3,245,750
Number of workers 48,146,471
Share of workers with any job change 0.504
SD of log wages 0.804
Summary of estimation results
SD of worker effects 0.638
SD of establishment effects 0.331
SD of covariates 0.159
SD of residuals 0.223
Correlation of worker and establishment effects 0.311
Adjusted $R^{2}$ of the model 0.895
Root Mean Squared Error of the model 0.261
Share of variance of log wages due to
Worker effects 0.629
Establishment effect 0.169
Covariance of worker and establishment effects 0.203
  • Note: The table summarizes the results from estimating the model $y_{it}=\alpha_{i}+\psi_{\bm{j}(i,t)}+X_{it}^{\prime}\beta+\varepsilon_{it}$, where the dependent variable is the log hourly wage of worker $i$ at year $t$. The model includes worker fixed effects $\alpha_{i}$ and establishment fixed effects $\psi_{\bm{j}(i,t)}$, as well as a set of covariates $X_{it}$ which contains year dummies and quadratic and cubic age, all interacted with dummies for five educational groups and with gender. The model is estimated on the largest connected set of establishments and workers aged 18-60 in RAIS between 2004 and 2008.
(1) (2) (3) (4) All candidates Candidates matched to RAIS Misconduct Elected Misconduct Elected Years of education -0.0017∗∗∗ 0.0045∗∗∗ -0.0014∗∗∗ 0.0051∗∗∗ (0.0000) (0.0001) (0.0001) (0.0001) Male -0.0057∗∗∗ 0.1009∗∗∗ -0.0012∗∗∗ 0.0715∗∗∗ (0.0004) (0.0006) (0.0006) (0.0009) Age 0.0001∗∗∗ -0.0004∗∗∗ 0.0002∗∗∗ -0.0013∗∗∗ (0.0000) (0.0000) (0.0000) (0.0001) Log wage -0.0025∗∗∗ 0.0377∗∗∗ (0.0004) (0.0007) AKM person effect -0.0029∗∗∗ 0.0364∗∗∗ (0.0003) (0.0005) AKM person effect residual -0.0016∗∗∗ 0.0230∗∗∗ (0.0002) (0.0005) Mean dep. variable 0.0358 0.1181 0.0313 0.1242 Observations 980,130 980,130 454,816 454,816 Note: The table shows correlations between various characteristics of councilor candidates and indicators for engaging in misconduct behavior during candidacy (columns (1) and (3)) and for being elected (columns (2) and (4)). The misbehavior variable captures irregularities detected by TSE during the elections in 2016 and 2020 that resulted in candidacy annulment, including abuse of power, failure to meet registration requirements, vote buying, prohibited conduct, ineligibility under the “Clean Record Act” (Lei da Ficha Limpa), illicit use of campaign resources, objections to the candidacy, rejection by the party or coalition, and invalidation of the party. See Section 6.1 for details on the measurements of the various candidate characteristics. In columns (1) and (2), the sample consists of all individuals who run for local councilor in the election years 2016 and 2020. Since TSE reports the socio-demographic information of candidates, we can show their correlation with misconduct and election outcomes for the full population of councilor candidates. For the wage measures shown in columns (3) and (4), we are limited to the subsample of candidates who appear in RAIS before the layoff period, i.e., those with at least one formal work contract between 2004 and 2008 (about 46% of all candidates).. Each cell shows the result of a bivariate OLS regression. Robust standard errors are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table D.2: Correlation of candidate characteristics with misconduct during candidacy and electoral performance
(1) (2) (3) (4) Outcome: $\Delta$ Share male (ppt) of … Councilor candidates Elected councilors [A] OLS $\Delta$ Employment rate (ppt) 0.078∗∗∗ 0.061∗∗∗ 0.100∗∗ 0.063 (0.020) (0.020) (0.050) (0.051) Observations 16,650 16,650 16,650 16,650 [B] 2SLS (Bartik IV) $\Delta$ Employment rate (ppt) 0.631∗∗∗ 0.631∗∗∗ 0.370∗∗ 0.354 (0.071) (0.082) (0.159) (0.189) KP F-stat 301.7 250.3 301.7 250.3 Observations 16,650 16,650 16,650 16,650 Fixed effects Year State $\times$ year Year State $\times$ year Note: The table presents OLS and 2SLS results from model (5) for the effect of formal employment rates on the male share of local councilors. The sample contains all municipal elections in 2004, 2008, 2012, and 2016. The dependent variable measures the percentage point change across subsequent elections in the male share of individuals running (columns (1) and (2)) and being elected (columns (3) and (4)) for local council. The explanatory variable is the percentage point change in the municipal formal employment-to-population rate. For the 2SLS results, changes in employment rates are instrumented with Bartik instruments that are constructed by interacting baseline differences in municipalities’ industry composition with national industry employment growth rates (see equation (6)). All regressions are weighted by municipal population. Standard errors clustered at the municipality level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table D.3: Municipality-level labor demand and gender of councilor candidates
(1) (2) (3) (4) Outcome: $\Delta$ Average age of … Councilor candidates Elected councilors [A] OLS $\Delta$ Employment rate (ppt) 0.022∗∗∗ 0.001 0.037∗∗ 0.027 (0.008) (0.009) (0.017) (0.018) Observations 16,650 16,650 16,650 16,650 [B] 2SLS (Bartik IV) $\Delta$ Employment rate (ppt) 0.081∗∗∗ 0.020 0.115∗∗ 0.111 (0.027) (0.032) (0.053) (0.062) KP F-stat 301.7 250.3 301.7 250.3 Observations 16,650 16,650 16,650 16,650 Fixed effects Year State $\times$ year Year State $\times$ year Note: The table presents OLS and 2SLS results from model (5) for the effect of formal employment rates on the average age of local councilors. The sample contains all municipal elections in 2004, 2008, 2012, and 2016. The dependent variable measures the change across subsequent elections in the average age of all individuals running (columns (1) and (2)) and being elected (columns (3) and (4)) for local council. The explanatory variable is the percentage point change in the municipal formal employment-to-population rate. For the 2SLS results, changes in employment rates are instrumented with Bartik instruments that are constructed by interacting baseline differences in municipalities’ industry composition with national industry employment growth rates (see equation (6)). All regressions are weighted by municipal population. Standard errors clustered at the municipality level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table D.4: Municipality-level labor demand and age of councilor candidates